OpenStudy (mayankdevnani):

CHALLENGING QUESTION !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Juliet has attempted 213 problems on Open Study and solved 210 of them correctly. Her friend Romeo has just joined Open Study, and has attempted 4 problems and solved 2 correctly. From now on, Juliet and Romeo will attempt all the same new problems. Find the minimum number of problems they must attempt before it is possible that Romeo's ratio of correct solutions to attempted problems will be greater than Juliet's.

mathslover (mathslover):

http://olimpiade.org/Forum/?qa=7848/brilliant-problem-solving Check that out please.

OpenStudy (mayankdevnani):

There is no exact answer

OpenStudy (mayankdevnani):

@hartnn

mathslover (mathslover):

You have the equation : 424 = n(5+n) You just need to solve this!

OpenStudy (anonymous):



OpenStudy (mayankdevnani):

by the way,what is the value of n ? @mathslover

OpenStudy (mayankdevnani):

@ehuman what is the final answer?

OpenStudy (anonymous):

on the 415th question Romeo can take the lead.

OpenStudy (mayankdevnani):

so 415 is the answer

OpenStudy (anonymous):

are you asking me or telling me?

OpenStudy (mayankdevnani):

asking

OpenStudy (anonymous):

they math is there in my drawing, unless I made an error in the initial formula, the answer is correct.

OpenStudy (mayankdevnani):

this is wrong answer

OpenStudy (mayankdevnani):

@Directrix

OpenStudy (anonymous):

ok reevaluating starting formula

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

@Hero

mathslover (mathslover):

Is it 414?

OpenStudy (mayankdevnani):

why?

mathslover (mathslover):

Yeah it seems to be ... I am getting the value of the variable that I had put there as negative.

OpenStudy (mayankdevnani):

HELP ME!!

OpenStudy (anonymous):

if you allow romeos average to start at 2/4 and juliets to start at 210/213, then in 415 100% correct answers later, romeo's average of .99760760555 will pass juliet's of .9952229299. The answer is 415

OpenStudy (anonymous):

just one question prior to that they both have the same average of .995215311

OpenStudy (mayankdevnani):

is he right? @mathslover and @Hero

mathslover (mathslover):

Well, I have tried 3-4 methods, one answer among them is 415 :D

mathslover (mathslover):

I have many varieties of answers :P

OpenStudy (mayankdevnani):

i don't know what is the answer

mathslover (mathslover):

Then I will go for 415

OpenStudy (mayankdevnani):

tell me your varieties @mathslover

OpenStudy (anonymous):

look at the formula in the drawing i did and point out what is in error with the starting formula, else it is indeed correct.

mathslover (mathslover):

415 should be correct one.

OpenStudy (anonymous):

414 = same ratio 415 = romeo has a better ratio

OpenStudy (mayankdevnani):

thank you guys ......for giving such a nice answer but i am still confused I'll be back day after tomorrow with a confirm answer. Again thank you guys . @Hero , @ehuman and @mathslover and please don't feel anything about me . If i hurt you, i am sorry. :)

OpenStudy (anonymous):

i'm pretty convinced that 415 is the answer

OpenStudy (anonymous):

nobody's hurt, just confident until proven otherwise :) @mayankdevnani

OpenStudy (anonymous):

\[Juliet: \frac{ 210+415 }{ 213+415 } = 0.995222929\]\[Romeo: \frac{ 2+415 }{ 4+415 } = 0.995226730\]

OpenStudy (anonymous):

plugging 414 yields the exact same answer (because we made it so)

hero (hero):

These numbers assume that Julie doesn't get any of the remaining 202 problems right.

hero (hero):

Which is ludicrous

OpenStudy (anonymous):

That isn't the case for the answer I gave. But you just made me realize that we're supposed to assume that she gets none right and that he gets them all right. "what is the MINIMUM amt of problems they must attempt" so 415 is not the answer. it is a smaller number

OpenStudy (anonymous):

\[\frac{ 210 }{ 213 +x } = \frac{ 2 + x }{ 4 + x }\]

hero (hero):

Actually \[\frac{ 210 }{ 213 +x } <\frac{ 2 + x }{ 4 + x }\]

OpenStudy (anonymous):

wow. turns out x > 18 only

OpenStudy (anonymous):

@mayankdevnani i hope you see this ^_^

OpenStudy (anonymous):

ohhhhh the minimum number, here I solved for the maximum number

mathslover (mathslover):

Nicely done @Euler271

OpenStudy (mayankdevnani):

so what is the answer? @Euler271

OpenStudy (anonymous):

x > 18